OpenStudy (anonymous):

solve : 2x over 3 - 4 over x= 5 over 3

OpenStudy (anonymous):

\[\frac{2x}{3} - \frac{4}{x} = \frac{5}{3}\] is this what the problem looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is x=4 and -1.5

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

\[\frac{2x(x)}{3(x)}-\frac{4(3)}{x(3)}=\frac{5}{3}\] \[\frac{2x ^{2}-12}{3x}=\frac{5}{3}\] multiply both sides by 3 \[\frac{2x ^{2}-12}{x}=5\] multiply both sides by x\[2x ^{2}-12=5x\] \[2x ^{2}-5x-12=0\] solve using the quadratic formula

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you're welcome